if f(x)+ square root x-5 and g(x) = -x find f+g and state the doamin
Answer by stanbon(75887)

if f(x)= square root x-5 and g(x) = -x find f+g and state the domain
-----------------
f(x) = sqrt(x-5) and g(x) = -x
---
(f+g)(x) = f(x)+g(x) = sqrt(x-5)+(-x)
-------------------
Domain is all Real Numbers >=5.
==================================
Cheers,
Stan H.